# Higher Order Structures For Graph Explanations

### v5

@@ 8 Conclusion and Future Work
We introduce FORGE, a novel framework that enhances GNN explainability by leveraging cell complexes. Our framework employs a novel lifting algorithm to convert graphs to cell complexes. Furthermore, we introduce information propagation algorithms to create more interpretable internal data representations. Our extensive evaluations demonstrate that FORGE consistently enhances explanation accuracy and faithfulness across both real-world and synthetic datasets. Our information propagation algorithms offer a versatile approach that can be extended to address various tasks involving higher-order structures. Future work could explore adaptive propagation approaches and investigate FORGE’s applicability to diverse graph learning tasks. We hope this work motivates further research into applying higher-order structures for enhanced interpretability.

### v6

@@ 8 Conclusion and Future Work
We introduce FORGE, a novel framework that enhances GNN explainability by leveraging cell complexes. Our framework employs a novel lifting algorithm to convert graphs to cell complexes. Furthermore, we introduce information propagation algorithms to create more interpretable internal data representations. Our extensive evaluations demonstrate that FORGE consistently enhances explanation accuracy and faithfulness across both real-world and synthetic datasets. Future work could explore adaptive propagation approaches and investigate FORGE’s applicability to diverse graph learning tasks. We hope this work motivates further research into applying higher-order structures for enhanced interpretability.
Acknowledgments
We are deeply grateful to the Precog research group for their collective support throughout this work. We extend special thanks to Shashwat Singh and Vamshi Krishna Bonagiri for their invaluable insights.